OpenStudy (anonymous):

the function y=f(x) is y=ax^3+2x^2 +bx. if the function has a max T.P at (3,36)..determine a and b values

OpenStudy (cwrw238):

first find f'(x) can you do that?

OpenStudy (cwrw238):

first find f'(x) can you do that?

OpenStudy (anonymous):

yes..so...3ax^2+4x+b

OpenStudy (cwrw238):

right this = 0 at a TP so 3a(3(^2 + 12 + b = 0 27a + b + 12 = 0

OpenStudy (cwrw238):

gotta go for a minute

OpenStudy (cwrw238):

gotta go for a minute

OpenStudy (anonymous):

hmm...so i use elimination or substitution method to solve for a and b?

OpenStudy (cwrw238):

you need another equation in a and b substitute y=36 and x = 3 into f(x): 36 = (3^3) a + 2(3)^2 + 3b

OpenStudy (cwrw238):

you need another equation in a and b substitute y=36 and x = 3 into f(x): 36 = (3^3) a + 2(3)^2 + 3b

OpenStudy (anonymous):

ok....thanks..i get it now

OpenStudy (cwrw238):

yw